OpenStudy (anonymous):

why does 3cos2x-2sin(x/3) never eqqual 5?

OpenStudy (science0229):

Hello

OpenStudy (science0229):

If that is true, than 3cos2x-2sin(x/3)=5 has no solutions, right?

OpenStudy (science0229):

First, we can put everything in terms of sin by using the identity cos2x=1-2(sinx)^2

OpenStudy (science0229):

3(1-2(sinx)^2)-2sin(x/3)=5

OpenStudy (science0229):

3-6(sinx)^2-2sin(x/3)=5 3(sinx)^2+sin(x/3)=-1

OpenStudy (science0229):

Now, this gets a bit interesting here.

OpenStudy (science0229):

What is the minimum of (sinx)^2?

OpenStudy (science0229):

@KerriA

OpenStudy (science0229):

Hello? Are you there?